From nch@roe.ac.uk Thu May 14 11:01:31 2009 Date: Fri, 8 May 2009 16:09:11 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Mark Holliman , Nigel Hambly , Nicholas Cross , Rob Blake , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Jim Emerson , Malcolm Stewart , Lorenzo Rimoldini , Mike Irwin , Peredur Williams , Bob Mann , Stephen Warren Subject: WFAU VDFS SCience Archive weekly project meeting minutes, 8/5/09 Minutes of WFAU VDFS Science Archive meeting: 8th May 2009 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, ETWS, NJC, MSH, MAR, RGM Apologies: JPE, JMS, AL, RSC, RPB, PMW DoNM: 10am, Friday May 15th 2009 in the Vista Hut Actions discharged this week: ----------------------------- None this week. Actions partly discharged but continuing: ----------------------------------------- ACTION: ETWS to parse that latest keyword info into the VSA metadata schema files. Continues; ETWS noted that communications with CASU indicate that the VIRCam data FITS keys are yet to be finalised, but that the procedure for parsing a standard set into an SQL schema is set up and can be tested further in the meantime. Actions carried forward from 1/05/09 meeting: ---------------------------------------------- None this week Specific points and new actions: -------------------------------- Project management: The nordic saga that is the Rolling Grant renewal continues. RGM noted the concern that STFC's Astronomy budget may fall short of being able to fund the WFAU RG at the originally envisaged level, so he and PMW are beginning to explore new avenues for future supplementary funding outwith the core roller. Apparently we're still at least a few weeks away from knowing where we stand with the funding level of the STFC renewal. WFCAM & VISTA updates: Good news from Hawaii: the weather has finally cleared up and lots of UKIDSS data has been obtained in the last week. Comments and issues arising from CASU fortnightly minutes: No new minutes this week. Networking: Transfers from CASU await release of 09A data; ETWS noted that our login to the CASU machines has stopped working, so we couldn't copy any data even if it had been released... has there been some kind of security audit at the CASU end, we wonder? WSA Operations: RPB has been concentrating on non-survey database preparation, with SW debugging/mods provided by RSC and NJC. Hardware: A new JBOD has arrived for the public server hosting the SDSS (and other external catalogue) DBs, so DR7 can be deployed. However RPB has been having trouble splitting the monolithic version supplied by Chicago (it has to be split otherwise it can't be hosted by our 32-bit Windows system). NCH has been looking into the problem with RPB. Software: ETWS noted: - Worked more on LGR's photometry code to update it for the latest software release. RSC reported: "I've been doing some more wider reviews of the software to see the best way to adapt it to work for both WSA & VSA curation. Noticed some minor issues along the way in the existing code that ETWS has helped clear up. The non-survey releases continue to test the software to its limits, so NJC and I have been coordinating fixes when RPB hits bugs. Also these releases have shown that the SQL Server parallelism bug is still causing us problems despite hyperthreading having been disabled on the load server, so I've been forced to apply more single-processor constraints on SQL queries generated by the CU6 script." NJC reported: "I have fixed some of the problems that affect curation of non-surveys, such as no matches for some sources when it came to calculating the astrometry statistics and the underlying cause of this in the best match table algorithm. I have also found a bug caused by updates to numpy which produced default values for the photometric statistics in the best match table and fixed a bug in the production of frame sets in the synoptic merge log table." Survey Data GPS survey head Phil Lucas has been in touch regarding GPS DR5. Apparently CASU will recatalogue the ~250 files with fluffy nebulosity by the end of the month. NCH suggested that we do the bulk of the DR5 source merging in advance of this to get it out of the way, then simply tidy up with a final run once the new catalogues are in. This needs the SW mods to CU7 for Phil's requested H2/K epoch-juggling to be implemented first - needs some thought by NCH and RSC. Non-survey Data Release: RPB has been working on further non-survey release DBs this week. Astrogrid deployment & Data Analysis services: MSH noted that he has been working on an application to transmogrify XML VOTable into KML, which is the (even more) verbose format used by GoogleSky applications to overplot catalogues in it's visualisations. Some discussion on how/where to deploy this, with the end result that it will be incoporated into the CEA suite, but also via some jsps linked from the main archive interactive interface pages to encourage those that are interested to play around with it. Miscellaneous: Nothing else this week. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. =============================================================